OpenStudy (anonymous):

find the value of:

OpenStudy (anonymous):

\[7x^2 \over \sqrt{x^5}\]

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {7x^2 \over \sqrt{x*x*x*x*x}} }\) \(\Large \color{Black}{\Rightarrow {7x^2 \over x^2\sqrt{x}} }\) \(\Large \color{Black}{\Rightarrow {7 \over \sqrt{x}} }\) \(\Large \color{Black}{\Rightarrow {7\sqrt{x} \over x} }\)